# iSeg: Interactive 3D Segmentation via Interactive Attention

### v1

@@ Appendix A Additional Results
Applications. In Fig. 11 in the main body, we show how our method’s results are used for local shape editing. iSeg selection corresponds to a contiguous region of an entire part of the shape. Thus, the segmented part can be easily manipulated. As the figure demonstrates, the entire bar of the chair is removed, only the head of the horse can be selected, the neck of the guitar is shortened, and a small local region of the candle’s flame is enlarged.
Limitation. Our method may not follow the symmetry of the mesh exactly, as exemplified in Fig. 19. For a click on the goat’s head, the segmented regions from the sides of the head differ somewhat from each other. The other face side is occluded from the clicked point, thus iSeg is not trained to segment those regions the same and does not produce a symmetric mask for both sides.
Appendix B Perceptual User Study
@@
The results from our perceptual study, reported in Tab. 1 in the paper, suggest that the segmented regions by iSeg are considered highly effective, compared to the partially effective segmentations by SAM baseline. This finding is consistent with the 2D nature of SAM. It is challenging to generate meaningful segmentation for occluded parts in 3D. For example, in Fig. 13 in the paper, the user’s point click is on one side of the camel’s hump. For a viewing angle where the click is not visible, the SAM baseline can only segment the hump partially. In contrast, iSeg fully segments the hump, even though the point click is concluded from the presented viewing angle. Due to iSeg’s 3D consistency, it can combine shape information from different viewing angles to produce an effective segmentation that is meaningful in 3D.
[…]
Figure 19. Limitation. iSeg may not produce a symmetric segmentation result for a symmetric shape. In this case, the segmented region of one side of the shape is different than the other.
Appendix C Ablation Test
@@
Appendix D Implementation Details
Encoder. We use the vertex coordinates of the mesh and apply positional encoding of size 515, which is the input to our encoder. The encoder is implemented as a multi-layer perceptron (mlp) of 6 layers, with 256 neurons per layer. Each layer except the last includes ReLU activation and layer normalization. After the last layer, we apply hyperbolic tangent activation without normalization. The encoder’s output is a per-vertex feature vector of size $256$.
Decoder. The interactive attention module contains linear layers of size $256\times 256$ for $W^{Q}$, $W_{pos}^{K}$, $W_{neg}^{K}$, $W_{pos}^{V}$, and $W_{neg}^{V}$. The output of this layer of size $256$ is concatenated to the per-vertex encoded feature vector of size $256$ to a total size of $512$. Then, a 16-layer mlp operates on the concatenated feature vector. The mlp has the first layer of size $512$, 14 layers of size $256$, and the last layer of size $2$. Each layer except the last includes ReLU activation and layer normalization. After the last layer, we apply a Softmax operation without normalization.
Compared method. As mentioned in the main body (Sec. 4.2), we compared the effectiveness of our method with the recent interactive segmentation work InterObject3D (Kontogianni et al., 2023). We adopted their publicly available pretrained model [^1] https://github.com/theodorakontogianni/InterObject3D/tree/main and added channels for point clicks, as instructed in their paper. For our experiments, we set the volume length for the point click to 0.05.
@@ Appendix D Implementation Details
Figure 21. Segmentation separability. The per-vertex segmentation probability can be separated distinctively into low and high value populations, which enables the hard selection of the segmented part by simple thresholding.

### v2

@@ B.2. Evaluaiton Metrics
B.3. SAM Baselines
As explained in Sec. 4.2 in the paper, we utilized the 2D segmentation model SAM (Kirillov et al., 2023) for segmenting shape parts. We evaluated several approaches to aggregate SAM’s predictions into 3D. In the first method, we applied a union between SAM’s masks (union aggregation). Meaning, a vertex belonged to the segmented region if it was segmented in one of the projected views of the shape. In the second method, we divided the number of times a vertex was segmented in the rendered views of the shape by the total number of rendered views (uniform average aggregation). In the third, instead of having the same division factor for all the vertices, we divided by the number of times a vertex was visible among the shape views (visibility average aggregation). Tab. 3 presents the segmentation performance for our quantitative evaluation on PartNet, as described in Secs. 4.2, B.1 and B.2. The visibility average aggregation had better results, thus, we used this baseline in the paper for comparison with our method.
Table 3. Quantitative PartNet results for baselines using SAM. We experimented with different methods of utilizing SAM for 3D part segmentation for shapes from the PartNet dataset. Considering SAM prediction averaged according to the vertex visibility (Visibility Average) yielded better results than the other options. Further details are given in Sec. B.3.
Aggregation Uniform Average Union Visibility Average
Accuracy $\uparrow$ 0.70 0.73 0.76
IoU $\uparrow$ 0.38 0.50 0.51
Appendix C Perceptual User Study
@@
Fig. 25 presents several example questions from our perceptual user study, demonstrating the comparison between the alternative techniques. The study’s results, reported in Tab. 2 in the paper, suggest that the segmented regions by iSeg are considered highly effective, compared to the partially effective segmentations by SAM baseline. This finding is consistent with the 2D nature of SAM. It is challenging to generate meaningful segmentation for occluded parts in 3D.
For example, in Fig. 15 in the paper, the user’s point click is on one side of the camel’s hump. For a viewing angle where the click is not visible, the SAM baseline can only segment the hump partially. In contrast, iSeg fully segments the hump, even though the point click is concluded from the presented viewing angle. Due to iSeg’s 3D consistency, it can combine shape information from different viewing angles to produce an effective segmentation that is meaningful in 3D.
Appendix D Ablation Test
@@
Appendix E Implementation Details
Encoder
We use the vertex coordinates of the mesh and apply positional encoding of size 515, which is the input to our encoder. The encoder is implemented as a multi-layer perceptron (MLP) of 6 layers, with 256 neurons per layer. Each layer except the last includes ReLU activation and layer normalization. After the last layer, we apply hyperbolic tangent activation without normalization. The encoder’s output is a per-vertex feature vector of size $256$. We trained the encoder with a wide range of random views from the full range $[-\frac{\pi}{2},\frac{\pi}{2})$ for the elevation angle and $[0,2\pi)$ for the azimuth angle. The resolution of the rendered image was $224\times 224$ pixels to match the resolution of the SAM ViT-H model used in our experiments.
Decoder
The interactive attention module contains linear layers of size $256\times 256$ for $W^{Q}$, $W_{pos}^{K}$, $W_{neg}^{K}$, $W_{pos}^{V}$, and $W_{neg}^{V}$. The output of this layer of size $256$ is concatenated to the per-vertex encoded feature vector of size $256$ to a total size of $512$. Then, a 16-layer MLP operates on the concatenated feature vector. The MLP has the first layer of size $512$, 14 layers of size $256$, and the last layer of size $2$. Each layer except the last includes ReLU activation and layer normalization. After the last layer, we apply a Softmax operation without normalization. As for the encoder, we used a wide range of random views for training the decoder (where the clicked points are visible), with the same $224\times 224$ image resolution.
[…]
Figure 23. Additional comparison results on the PartNet dataset. Each pair shows different views of the segmentation result for different interactive techniques. iSeg produces better segmentation results, than the compacted methods, yielding segmentations that are close to the ground-truth part annotations.
Compared method
As mentioned in the main body (Sec. 4.2), we compared the effectiveness of our method with the recent interactive segmentation work InterObject3D (Kontogianni et al., 2023). Since the training code was unavailable, we adopted their publicly available pretrained model [^1] https://github.com/theodorakontogianni/InterObject3D/tree/main and used channels for point clicks, as instructed in their paper. For our experiments, we set the volume length for the point click to be 0.05.
@@ Appendix E Implementation Details
Figure 25. Perceptual study example questions. For each mesh, we ask the participants to rate the effectiveness of the segmentation for the compared methods, where they are unaware of which method is used. The other techniques are perceived as partially effective, as they do not mark the entire 3D region corresponding to the clicked point. In contrast, iSeg selects complete regions in a 3D-consistent fashion and is considered highly effective.